A tennis player serves a tennis ball such that it is moving horizontally when it leaves the racquet. When the ball travels a horizontal distance of 16 m, it has dropped 59 cm from its original height when it left the racquet. What was the initial speed, in m/s, of the tennis ball? (Neglect air resistance.)
Answer #1

From vertical motion , y = 0.59m , a = 9.8m/s^2 and initial velocity It = 0

=) time t = (2y/a)^1/2 = ( 2×0.59/9.8)^0.5 = 0.347s

Now from horizontal motion , x = 16m and t = 0.347s

Hence Ux = 16/0.347 = 46.11m/s